# Michael Callen Miguel Fajardo-Steinhäuser Michael G. Findley

### v2

@@ 2.3 Intervention & Randomization
Uplift Afghanistan, a nonprofit organization based in the U.S., transferred digital value vouchers to beneficiaries’ mobile phones that were redeemable at HesabPay’s acceptance network of local merchants for any available goods. Specifically, the intervention transferred 4,000 AFN (approximately \$45 USD) every two weeks for two months to households in 16 urban neighborhoods in Kabul, Herat, and Mazar-i-Sharif. From our experimental sample of 2,409 households, we randomly assigned 1,208 households to an “early" group, which received benefits from November 6, 2022 until December 31, 2022 (henceforth, treatment group). The remaining 1,201 households formed the “late" group, which received benefits from January 1, 2023 (two weeks after the “early" group stopped receiving payments) until February 28, 2023 (henceforth, control group). The randomization was stratified on two variables: The nahia (neighborhood) in which they registered (our study includes beneficiaries from neighborhoods), and a measure of vulnerability. As can be observed in our PAP, the treatment groups are balanced on 17 of the 18 outcome and heterogeneity variables we collected at baseline. [^13] As we note on P. 15 of the PAP, with 18 balance variables and a 5% significance level, “we would expect to see p-values of less than 0.05 in 2 of the variables.” We thus committed to implement the first randomization draw with less than two variables with a p-value below 0.05. This condition was met on the first random seed we tried, resulting in one unbalanced variable out of 18.
@@ 2.4 Data collection
Data were collected four ways. First, a baseline survey was completed during the onboarding session (see above). Second, we had access to participants’ transaction data from the mobile payments provider, which we could link to participants’ survey data. We obtained permission to do so during the consent process. Third, we conducted four rounds of follow-up surveys over the phone. [^16] All surveys were conducted by female enumerators, consistent with local norms. Participants were informed that the survey was completely voluntary and would not affect their aid payments, and they could skip any question they did not want to answer. Participants received a 350 AFA ($\sim$4 USD) payment for completing each survey. Overall, we have response rates of about 99% across survey rounds, with no difference in response rates between the treatment and control groups. Fourth, we attempted to survey of all 26 merchants who facilitated initial test transactions with beneficiaries, and succeeded in contacting 19, to confirm how beneficiaries had used their payments and to check whether they were asked to provide favors or extra-legal taxes to local authorities.
We pre-specified all analyses in a PAP, including how the outcome variables would be constructed, and what our primary outcomes variables were going to be. We divided our outcome variables in three families: Basic needs (also called “food security" in the manuscript), well-being and informal taxation (also called “diversion" in the manuscript). In doing so, we tried to provide a level of detail consistent with the standard articulated in (Duflo et al., 2020). This standard requires that two research assistants can take the data and the PAP, and with only these two items separately produce identical analysis. For each of these three families of outcomes, we create a summary index following (Kling, Liebman, and Katz, 2007).
@@ 2.5 Methodology
To estimate treatment effects, we use linear regressions (see Online Appendix 2 for estimating equations). We study the intervention’s effect by regressing the pre-specified outcome variables on an indicator for treatment assignment (baseline analysis). We control for the outcome variable’s baseline value (if available), strata fixed effects, and survey round fixed effects. Standard errors are clustered at the level of treatment assignment (the participant). Note that due to very high response rates and no compliance issues, we do not conduct treatment-on-the-treated regressions. To avoid false discoveries due to multiple hypothesis testing, we control for the Family Wise Error Rate (FWER) for each family of primary outcomes to be conservative, and control for the False Discovery Rate for secondary outcomes as pre-specified in our PAP.
We conduct several complementary analyses as well, which were also pre-specified. First, we study how the effects change over time by including an interaction term between the treatment assignment indicator and an indicator for the second survey round. We also study dynamic effects graphically by computing the means of given outcome variables at the survey week level for treatment and control groups separately, as shown in Figure A2. Second, we study whether the effects are heterogeneous along several (pre-specified) variables, by interacting each of these variables with the treatment assignment indicator.
@@ 3.1 Can extremely poor women use digital transfers?
Beyond the high levels of usage, three additional results support the argument that this tech-illiterate population can use digital transfers. First, 20.9% of the funds were spent at different merchants than those who facilitated an initial test transaction during onboarding, indicating that beneficiaries understood they could use the payment at any participating merchant. Second, following our PAP, we also checked whether impacts on food security outcomes vary by pre-intervention need, city, marital status, age, household size, education, and whether the recipient was the primary household financial decision-maker. We found no heterogeneity in impacts, consistent with the technology being roughly equally useful for the different groups in our study (Figure A3).
Finally, we asked participants in a hypothetical exercise whether they would prefer to receive their full 4000 AFN payments digitally, or $4000-X$ AFN in cash, where $X$ was either 100 AFN, 300 AFN, 500 AFN. These amounts might reflect the costs required to provide aid as physical cash (see Online Appendix 2 for details). With a conservative fee of 100 AFN (2.5%), 80% of participants preferred digital aid over cash; with a higher fee of 300 AFN (7.5%), the share choosing digital aid over cash increased to $\sim$95(Figure A4). Collectively, the results indicate that digital payments are a viable option even when people have limited experience using digital technology.
@@ 3.2 Do digital transfers improve food security and mental well-being?
The payments reduced all four pre-specified measures of food security by meaningful amounts (see Table 1, Panel A for estimates and control means). Beneficiaries, on average, skipped meals in 0.76 fewer days per week (SE = 0.051, $p$leq 0.0001), children were 11.7 percentage points less likely to have skipped meals over the past week (SE = 0.012, $p$leq 0.0001), the share of households where everyone was able to eat regularly during the prior week increased by 9.3 percentage points (SE = 0.015, $p$leq 0.0001), and beneficiaries reduced meals of only bread and tea by 1.608 (SE = 0.121, $p$leq 0.0001). An index of these four measures – constructed as the average of the standardized measures following (Kling, Liebman, and Katz, 2007) – improved by 0.5 SDs (SE = 0.032, $p$leq 0.0001).
@@
Under secondary outcomes, we find evidence of a more diverse diet (see Table 1, Panel B for estimates and control means). With a recall period of one week prior to the survey, beneficiaries ate rice on 0.6 days (SE = 0.035, $p$leq 0.0001), beans on 0.49 days (SE = 0.029, $p$leq 0.0001), chicken on 0.01 days (SE = 0.016, $p$ = 0.035), and dairy on 0.05 days (SE = 0.013, $p$leq 0.0001) more than the control group. Participants also report an increased ability to purchase medicine when needed. Consistent with pre-registered expectations in our PAP, we do not find any increases in outside income, employment, or agency over financial decisions. [^19] Regarding income and employment, we write in our PAP: “due to the existing restrictions on women’s liberties in Afghanistan, we believe it is unlikely that [these] will change.”
@@
Primary food security outcomes kept improving for the two months during which beneficiaries were paid (Table 2), and more modest improvements remain for at least two months after payments conclude (Figure A2). Regarding dietary diversity, estimates reported in Table A2 indicate that consumption of beans and rice is increasing over time.
@@
To assess if these estimated impacts are in line with what we should expect given the size of the payments, we collected data from a subsample of households on what they were purchasing (e.g., wheat flour, cooking oil, and sugar) and price data from merchants. We calculate that the cost of a basket of popular goods to support a family for two weeks is consistent with the aid payment size plus the survey participation incentive (Bete, 2022). While not eliminated, skipped meals declined after each payment (Figure A2, Panel A). Online Appendix 5 provides further discussion of the magnitudes of needs results.
@@
Turning to mental well-being, since the Taliban takeover in 2021, Afghanistan consistently ranks as the country with the lowest levels of happiness (Gallup, 2022a, c). Our participants report extremely low levels of happiness and financial health. Treatment improved all four prespecified measures of mental and financial health, with an index of these measures improving by 1.5 SDs (Table 1, Panel B), although from a very low base and these impacts disappear as soon as payment stops (Figure A2). Specifically, beneficiaries are 33.5 percentage points more likely to report that they feel the economic situation of their household has improved compared to 30 days ago (SE = 0.011, $p$leq 0.0001), from a base of just 4.8% in the control group. They are also 26.3 percentage points more likely to report being satisfied with their current financial situation (SE = 0.012, $p$leq 0.0001). The intervention also increased measures of mental well-being: beneficiaries are 28 percentage points more likely to report being very or quite happy (SE = 0.014, $p$leq 0.0001), and report a score 1.96 higher on the Cantril Self-Anchoring Striving Scale, an established measure of life satisfaction that runs from 1 to 10 (SE = 0.068, $p$leq 0.0001), relative to the control group. [^20] While these results may seem unrealistically large, this is attributable to low base levels, as is seen in other surveys in contemporary Afghanistan (Gallup, 2022a, c). These effects are increasing over time, with the effects after receiving 3-4 aid payments between 40% to 100% larger than after 1-2 aid payments (Table 2, Panel B). [^21] Our food security results complement a broader literature demonstrating mostly positive impacts, whereas our mental well-being results contribute newer insights to a more nascent literature. As discussed in Online Appendix 2, our review of 23 cash-based programs in humanitarian contexts found that 21 included food security outcomes, with 16 of those 21 finding any positive effect. Less attention has focused on the mental impacts of cash-based programs, with only seven of the 23 studies considering a relevant outcome and six finding any positive effect.
@@
Our results are robust to common concerns in Randomized Control Trials (RCTs). First, we find no evidence of treatment beneficiaries providing part of their payments to control beneficiaries or of control beneficiaries borrowing against payments they would receive in the last two months of the program (see Online Appendix 3). Second, survey bias in the absence of objective measures of well-being (such as anthropometric or biometric measures) is a concern, especially since participants might respond strategically if they believe this could influence their benefits. To check for potential strategic reporting by participants, we primed half of our respondents with an explicit description of the study’s purpose, finding no evidence that the prime influenced responses (see Online Appendix 3 and Table A3). Finally, given that participants had the option of skipping questions, and to keep the sample comparable, we re-estimate the results restricting the sample to those who answered all questions relevant for the analysis. Results using this restricted sample are almost identical to those estimated on the complete sample (see Table A4).
@@ 3.3 Are digital transfers diverted?
There are at least four ways that hostile regimes might capture aid. First, they might influence who is eligible, including by creating fictitious “ghost" beneficiaries. Second, they can ask recipients to hand over aid transfers. Third, they can capture aid while in transit. Fourth, they can ask merchants or payments platforms who are serving aid beneficiaries to pay bribes or additional tax. We organize the presentation of results around these four potential diversion strategies and discuss whether digital delivery might constrain them. Such diversion strategies are widely documented both in Afghanistan (Special Inspector General for Afghanistan Reconstruction, 2023b; Sopko, 2023) and in other countries (see Online Appendix 1 for a discussion of diversion of humanitarian aid in the literature).
@@
In our study, beneficiaries were identified by local elected community councils in consultation with communities using a “well-being analysis" described in Online Appendix B. The prevalence of mobile phones enables phone surveys as a means of checking that beneficiaries are indeed vulnerable, even at large-scale. Our surveys confirm that beneficiaries were quite vulnerable (see Table A5). The digital payments platform also requires Know Your Customer identity verification, which can automatically compare beneficiary names against sanctions lists to further guard against capture by hostile actors. Evidence from a diverse set of global contexts (Guggenheim and Petrie, 2022; Samii, 2023) and the pre-2021 Afghanistan NSP program specifically (Beath, Christia, and Enikolopov, 2018; Burde et al., 2023) also indicates that most diversion occurs at national and regional levels and primarily through bureaucratic and partisan channels. At the local level, in contrast, CDCs develop transparency and monitoring mechanisms to prevent diversion, making it more challenging for local authorities to interfere in influencing eligibility.
@@
To check for taxation, we directly asked beneficiaries whether they have been asked for informal assistance. Specifically, we asked them whether local community leaders or government officials have asked them for any kind of assistance, such as food or money. Given that participants may be reluctant to disclose doing this themselves, we first asked them whether they know someone in their community who has been asked to do so (Reinikka and Svensson, 2006), and then whether they themselves had been asked. Results are presented in the first four columns of Table 3, Panel A, which show that the treatment group does not report informal payments to authorities in larger proportion than the control group regardless of how the question was framed (all coefficients are insignificant at traditional levels, and precisely estimated). When looking at the KLK Index combining the four individual questions (column 5), the coefficient is marginally significant at the 10% level. However, this is driven by a single individual who answered yes to three of the four questions, which due to the way the index is constructed, receives an extremely high index value of 34 standard deviations. Removing this observation leads the results using the KLK Index to lose statistical significance at conventional levels (column 4 in Table A6).
@@
Importantly, once beneficiaries are onboarded, distribution does not require travel outside the home as it might with food or cash distributions. The distribution of digital aid is therefore less visible than cash or in-kind distribution. Finally, the distribution of digital payments is instantaneous, with recipients able to access their funds immediately. Figure 2, Panel A displays the share of the (cumulative) funds participants have received that remains unspent over time, and shows that participants spent over 70% of the funds they received the day they received their first payment. This drops quickly over the next few days to below 5% of the received funds. A similar pattern holds every time participants receive their funds. Panel B displays how participants spent their money on a daily basis.
@@
The government has the legal authority to tax both merchants and digital payment platforms. To check for merchant taxation, we surveyed all 26 merchants who assisted beneficiaries with an initial test transaction to ask about extra-legal taxes (merchants do need to pay ordinary sales tax on any transactions). None of the merchants reported paying extra-legal taxes, being asked to provide favors, or being aware of other merchants paying extra-legal taxes or providing favors.

### v4

@@ 2.5 Intervention & Randomization
Uplift Afghanistan transferred digital value vouchers to beneficiaries’ mobile phones that were redeemable at HesabPay’s acceptance network of local merchants for any available goods. Specifically, the intervention transferred 4,000 AFN (approximately \$45 USD) every two weeks for two months to households in 16 urban neighborhoods in Kabul, Herat, and Mazar-i-Sharif. From our experimental sample of 2,409 households, we randomly assigned 1,208 households to an “early” group, which received benefits from November 6, 2022 until December 31, 2022 (henceforth, treatment group). The remaining 1,201 households formed the “late” group, which received benefits from January 1, 2023 (two weeks after the “early” group stopped receiving payments) until February 28, 2023 (henceforth, control group). The randomization was stratified on two variables: The nahia (neighborhood) in which they registered (our study includes beneficiaries from neighborhoods), and a measure of vulnerability. [^15] Participants in the same nahia had access to the same merchant acceptance network, motivating this decision. We discuss the stratification procedure in detail in our pre-analysis plan. As can be observed in A2, the treatment groups are balanced on 17 of the 18 outcome and heterogeneity variables we collected at baseline. [^16] As we note on P. 15 of the PAP, with 18 balance variables and a 5% significance level, “we would expect to see p-values of less than 0.05 in 2 of the variables.” We thus committed to implement the first randomization draw with less than two variables with a p-value below 0.05. This condition was met on the first random seed we tried, resulting in one unbalanced variable out of 18.
@@ 2.6 Data collection
Data were collected four ways. First, a baseline survey was completed during the onboarding session (see above). Second, we had access to participants’ transaction data from the mobile payments provider, which we could link to participants’ survey data. We obtained permission to do so during the consent process. Third, we conducted four rounds of follow-up surveys over the phone. [^19] All surveys were conducted by female enumerators, consistent with local norms. Participants were informed that the survey was completely voluntary and would not affect their aid payments, and they could skip any question they did not want to answer. Participants received a 350 AFA ($\sim$4 USD) payment for completing each survey (see Online Appendix D). Overall, we have response rates of about 99% across survey rounds, with no difference in response rates between the treatment and control groups. Fourth, we attempted to survey all 26 merchants who facilitated initial test transactions with beneficiaries, and succeeded in contacting 19, to confirm how beneficiaries had used their payments and to check whether they were asked to provide favors or extra-legal taxes to local authorities.
We pre-specified all analyses in a PAP, including how the outcome variables would be constructed, and what our primary outcomes variables were going to be. We divided our outcome variables in three families: Basic needs (also called “food security” in the manuscript), well-being and informal taxation (also called “diversion” in the manuscript). In doing so, we tried to provide a level of detail consistent with the standard articulated in Duflo et al. (2020). This standard requires that two research assistants can take the data and the PAP, and with only these two items separately produce identical analysis. For each of these three families of outcomes, we create a summary index following Kling, Liebman, and Katz (2007).
@@ 2.7 Methodology
To estimate treatment effects, we use linear regressions (see Online Appendix 2 for estimating equations). We study the intervention’s effect by regressing the pre-specified outcome variables on an indicator for treatment assignment (baseline analysis). We control for the outcome variable’s baseline value (if available), strata fixed effects, and survey round fixed effects. Standard errors are clustered at the level of treatment assignment (the participant). Note that due to very high response rates and no compliance issues, we do not conduct treatment-on-the-treated regressions. To avoid false discoveries due to multiple hypothesis testing, we control for the Family Wise Error Rate (FWER) for each family of primary outcomes and control for the False Discovery Rate for secondary outcomes as pre-specified in our PAP.
We conduct several complementary analyses as well, which were also pre-specified. First, we study how the effects change over time by including an interaction term between the treatment assignment indicator and an indicator for the second survey round. We also study dynamic effects graphically by computing the means of given outcome variables at the survey week level for treatment and control groups separately, as shown in Figure 2. Second, we study whether the effects are heterogeneous along several (pre-specified) variables, by interacting each of these variables with the treatment assignment indicator.
@@ 3.1 Can extremely poor women use digital transfers?
Beyond the high levels of usage, three additional results support the argument that this tech-illiterate population can use digital transfers. First, 20.9% of the funds were spent at different merchants than those who facilitated an initial test transaction during onboarding, indicating that beneficiaries understood they could use the payment at any participating merchant. Second, following our PAP, we also checked whether impacts on food security outcomes vary by pre-intervention need, city, marital status, age, household size, education, and whether the recipient was the primary household financial decision-maker. We found no heterogeneity in impacts, consistent with the technology being roughly equally useful for the different groups in our study (A2).
Finally, we asked participants in a hypothetical exercise whether they would prefer to receive their full 4000 AFN payments digitally, or $4000-X$ AFN in cash, where $X$ was either 100 AFN, 300 AFN, 500 AFN. These amounts might reflect the costs required to provide aid as physical cash (see Online Appendix 2 for details). With a conservative fee of 100 AFN (2.5%), 80% of participants preferred digital aid over cash; with a higher fee of 300 AFN (7.5%), the share choosing digital aid over cash increased to $\sim$95(A3). Collectively, the results indicate that digital payments are a viable option even when people have limited experience using digital technology.
@@ 3.2 Do digital transfers improve food security and mental well-being?
The payments reduced all four pre-specified measures of food security by meaningful amounts (see Table 1, Panel A for estimates and control means). Beneficiaries, on average, skipped meals in 0.76 fewer days per week (SE = 0.051, $p$leq 0.0001), children were 11.7 percentage points less likely to have skipped meals over the past week (SE = 0.012, $p$leq 0.0001), the share of households where everyone was able to eat regularly during the prior week increased by 9.3 percentage points (SE = 0.015, $p$leq 0.0001), and beneficiaries reduced meals of only bread and tea by 1.608 (SE = 0.121, $p$leq 0.0001). An index of these four measures – constructed as the average of the standardized measures following Kling, Liebman, and Katz (2007) – improved by 0.5 SDs (SE = 0.032, $p$leq 0.0001).
@@
Under secondary outcomes, we find evidence of a more diverse diet (see Table 1, Panel B for estimates and control means). With a recall period of one week prior to the survey, beneficiaries ate rice on 0.6 days (SE = 0.035, $p$leq 0.0001), beans on 0.49 days (SE = 0.029, $p$leq 0.0001), chicken on 0.01 days (SE = 0.006, $p$ = 0.035), and dairy on 0.05 days (SE = 0.013, $p$leq 0.0001) more than the control group. Participants also report an increased ability to purchase medicine when needed. Consistent with pre-registered expectations in our PAP, we do not find any increases in outside income, employment, or agency over financial decisions. [^22] Regarding income and employment, we write in our PAP: “due to the existing restrictions on women’s liberties in Afghanistan, we believe it is unlikely that [these] will change.”
@@
Primary food security outcomes kept improving for the two months during which beneficiaries were paid (Table 2), and more modest improvements remain for at least two months after payments conclude (Figure 2). Regarding dietary diversity, estimates reported in A3 indicate that consumption of beans and rice is increasing over time.
@@
To assess if these estimated impacts are in line with what we should expect given the size of the payments, we collected data from a subsample of households on what they were purchasing (e.g., wheat flour, cooking oil, and sugar) and price data from merchants. We calculate that the cost of a basket of popular goods to support a family for two weeks is consistent with the aid payment size plus the survey participation incentive (Bete, 2022). While not eliminated, skipped meals declined after each payment (Figure 2, Panel A). Online Appendix 5 provides further discussion of the magnitudes of needs results.
@@
Turning to mental well-being, since the Taliban takeover in 2021, Afghanistan consistently ranks as the country with the lowest levels of happiness (Gallup, 2022a, c). Our participants report extremely low levels of happiness and financial health. Treatment improved all four prespecified measures of mental and financial health, with an index of these measures improving by 1.5 SDs (Table 1, Panel B), although from a very low base and these impacts disappear as soon as payment stops (Figure 2). Specifically, beneficiaries are 33.5 percentage points more likely to report that they feel the economic situation of their household has improved compared to 30 days ago (SE = 0.011, $p$leq 0.0001), from a base of just 4.8% in the control group. They are also 26.3 percentage points more likely to report being satisfied with their current financial situation (SE = 0.012, $p$leq 0.0001). The intervention also increased measures of mental well-being: beneficiaries are 28 percentage points more likely to report being very or quite happy (SE = 0.014, $p$leq 0.0001), and report a score 1.96 higher on the Cantril Self-Anchoring Striving Scale, an established measure of life satisfaction that runs from 1 to 10 (SE = 0.068, $p$leq 0.0001), relative to the control group. [^23] While these results may seem unrealistically large, this is attributable to low base levels, as is seen in other surveys in contemporary Afghanistan (Gallup, 2022a, c). These effects are increasing over time, with the effects after receiving 3-4 aid payments between 40% to 100% larger than after 1-2 aid payments (Table 2, Panel B). [^24] Our food security results complement a broader literature demonstrating mostly positive impacts, whereas our mental well-being results contribute newer insights to a more nascent literature. As discussed in Online Appendix 2, our review of 23 cash-based programs in humanitarian contexts found that 21 included food security outcomes, with 16 of those 21 finding any positive effect. Less attention has focused on the mental impacts of cash-based programs, with only seven of the 23 studies considering a relevant outcome and six finding any positive effect.
@@
Our results are robust to common concerns in Randomized Control Trials (RCTs). First, we find no evidence of treatment beneficiaries providing part of their payments to control beneficiaries or of control beneficiaries borrowing against payments they would receive in the last two months of the program (see Online Appendix 3). Second, given that participants had the option of skipping questions, and to keep the sample comparable, we re-estimate the results restricting the sample to those who answered all questions relevant for the analysis. Results using this restricted sample are almost identical to those estimated on the complete sample (see A5).
Finally, survey bias in the absence of objective measures of well-being (such as anthropometric or biometric measures) is a concern, especially since participants might respond strategically if they believe this could influence their benefits. To check for potential strategic reporting by participants, we primed half of our respondents with an explicit description of the study’s purpose. Specifically, in the second round of follow-up surveys, we randomly assigned individuals into two groups: a “primed” group hears the following statement just before the questions related to needs: “I would now like to ask you a few questions about how you and your family are doing. The goal of the CDDO and HesabPay program is to help you and your family meet basic needs, such as buying food, and we would like to see how you are doing in this regard. We will share what we learn from interviewing participants like yourself, with international organizations who are trying to help Afghans deal with these difficult times.” Thus, this group is explicitly told what we are expecting to find. The “not primed” group hears this placebo statement instead: “I would now like to ask you a few questions about how you and your family are doing.” We find no evidence that the prime influenced responses (see Online Appendix 3 and A4).
@@ 3.3 Are digital transfers diverted?
There are at least four ways that hostile regimes might capture aid. First, they might influence who is eligible, including by creating fictitious “ghost” beneficiaries. Second, they can ask recipients to hand over aid transfers. Third, they can capture aid while in transit. Fourth, they can ask merchants or payments platforms who are serving aid beneficiaries to pay bribes or additional tax. We organize the presentation of results around these four potential diversion strategies and discuss whether digital delivery might constrain them. Such diversion strategies are widely documented both in Afghanistan (Special Inspector General for Afghanistan Reconstruction, 2023b; Sopko, 2023) and in other countries (see Online Appendix 1 for a discussion of diversion of humanitarian aid in the literature).
@@
In our study, beneficiaries were identified by local elected community councils in consultation with communities using a “well-being analysis” described in Online Appendix B. The prevalence of mobile phones enables phone surveys as a means of checking that beneficiaries are indeed vulnerable, even at large-scale. Our surveys confirm that beneficiaries were quite vulnerable (see A6). The digital payments platform also requires Know Your Customer identity verification, which can automatically compare beneficiary names against sanctions lists to further guard against capture by hostile actors. Evidence from a diverse set of global contexts (Guggenheim and Petrie, 2022; Samii, 2023) and the pre-2021 Afghanistan NSP program specifically (Beath, Christia, and Enikolopov, 2018; Burde et al., 2023) also indicates that most diversion occurs at national and regional levels and primarily through bureaucratic and partisan channels. At the local level, in contrast, CDCs develop transparency and monitoring mechanisms to prevent diversion, making it more challenging for local authorities to interfere in influencing eligibility.
@@
To check for taxation, we directly asked beneficiaries whether they have been asked for informal assistance. Specifically, we asked them whether local community leaders or government officials have asked them for any kind of assistance, such as food or money. Given that participants may be reluctant to disclose doing this themselves, we first asked them whether they know someone in their community who has been asked to do so (Reinikka and Svensson, 2006), and then whether they themselves had been asked. Results are presented in the first four columns of Table 3, Panel A, which show that the treatment group does not report informal payments to authorities in larger proportion than the control group regardless of how the question was framed (all coefficients are insignificant at traditional levels, and precisely estimated). When looking at the KLK Index combining the four individual questions (column 5), the coefficient is marginally significant at the 10% level. However, this is driven by a single individual who answered yes to three of the four questions, which due to the way the index is constructed, receives an extremely high index value of 34 standard deviations. Removing this observation leads the results using the KLK Index to lose statistical significance at conventional levels (column 4 in A7).
@@
Importantly, once beneficiaries are onboarded, distribution does not require travel outside the home as it might with food or cash distributions. The distribution of digital aid is therefore less visible than cash or in-kind distribution. Finally, the distribution of digital payments is instantaneous, with recipients able to access their funds immediately. Figure 3, Panel A displays the share of the cumulative funds participants have received that remains unspent over time, and shows that participants spent over 70% of the funds they received the day they received their first payment. This drops quickly over the next few days to below 5% of the received funds. A similar pattern holds every time participants receive their funds. Panel B displays how participants spent their money on a daily basis.
@@
The government has the legal authority to tax both merchants and digital payment platforms. To check for merchant taxation, we surveyed 26 merchants who assisted beneficiaries with an initial test transaction to ask about extra-legal taxes (merchants do need to pay ordinary sales tax on any transactions). None of the merchants reported paying extra-legal taxes, being asked to provide favors, or being aware of other merchants paying extra-legal taxes or providing favors.